OpenStudy (watchmath):

Repost: Let \(x_1,\ldots,x_7\) be real numbers such that \[\begin{align*} x_1+4x_2+9x_3+16x_4+25x_5+36x_6+49x_7&=1\\ 4x_1+9x_2+16x_3+25x_4+36x_5+49x_6+64x_7&=12\\ 9x_1+16x_2+25x_3+36x_4+49x_5+64x_6+81x_7&=123 \end{align*}\] Compute \(16x_1+25x_2+36x_3+49x_4+64x_5+81x_6+100x_7\)

OpenStudy (dumbcow):

\[\sum_{n=1}^{7}n ^{2}X _{n}=1\] \[\sum_{n=1}^{7}(n+1)^{2}X _{n} = \sum_{n=1}^{7}n^{2}X _{n}+2nX _{n}+X _{n}=12\] \[\sum_{n=1}^{7}(n+2)^{2}X _{n} = \sum_{n=1}^{7}n^{2}X _{n}+4nX _{n}+4X _{n}=123\] \[u = \sum_{n=1}^{7}nX _{n}\] \[v = \sum_{n=1}^{7}X _{n}\] System of 2 equations \[1+2u + v = 12\] \[1+4u+4v = 123\] solving yields \[u = -\frac{39}{2}\] \[v=50\] \[\sum_{n=1}^{7}(n+3)^{2}X_n=1+6u+9v =334\]

OpenStudy (watchmath):

Here is another way \(n^2-3(n+1)^2+3(n+2)^2=(n+3)^2\) for all \(n\) It follows that \(16x_1+25x_2+36x_3+49x_4+64x_5+81x_6+100x_7=1-3(12)+3(123)=334\)

OpenStudy (dumbcow):

yes less work too never thought of that nice :)